OpenStudy (highschoolmom2010):

What is the equation of y = x^3 with the given transformations? horizontal translation right 10 units and vertical translation down 12 units

OpenStudy (anonymous):

here are the rules

OpenStudy (anonymous):

f(x)= x^3 f(x)+a= x^3 +a up a units f(x-b)= (x-b)^3 right b units. Be so careful here cuzz it is - not +. Small quirk and finally you can put it all together

OpenStudy (anonymous):

Plug in your numbers and put it all together. Let me know if you need help doing that.

OpenStudy (highschoolmom2010):

F(x)-12=x^3-12 F(x-10)=x^3-10 Like that

OpenStudy (anonymous):

can you please help me :)

OpenStudy (anonymous):

well let me give you a hint. To put it all together youd do this f(x-b)+a = (x-b)^3+a now plug in again :)

OpenStudy (highschoolmom2010):

F(x-10)+12=x-10^3+12

OpenStudy (anonymous):

you forgot perenthesis :) show me where :) also F(x-10)+12=x-10^3+12 ^ forget the ^3

OpenStudy (highschoolmom2010):

F(x-10)^3+12=(x-10)^3+12

OpenStudy (anonymous):

Yes i think so :D